# LiDAR 2.0: Hierarchical Curvy Waveguide Detailed Routing for Large-Scale Photonic Integrated Circuits

### v1

@@ VI Evaluation Results
The proposed photonic detailed routing framework is implemented in Python based on GDSFactory 8 [38] libraries. All evaluations are conducted on a on a Linux server with a 128-core AMD EPYC 7763 CPU and 512-GB memory. Note that runtime and results may slightly differ from the ISPD version due to platform and GDSFactory version variations.
@@ VI-A Results on Spacious Benchmarks
Analysis of PTC Results. Despite the Clements circuit having a regular topology with no inherent crossings and relatively spacious layouts, insufficient planning and lack of reserved access space lead to unnecessary crossings and DRVs in PROTON. These routing conflicts become more severe in ADEPT benchmarks, which features dense multi-port MMIs and a large number of crossings. By incorporating Accessibility-Enhanced Port Assignment and Port-Group-Based Net Ordering, LiDAR 2.0 effectively mitigates these conflicts and reduces unnecessary rip-up and reroute. As circuit scale increases, PROTON suffers from escalating #DRVs and runtime, while LiDAR 2.0 consistently produces DRV-free, low-IL layouts.
Analysis of Optical Switch Results. Light benchmark provides abundant routing resources and is primarily used to evaluate the router’s ability to optimize insertion loss and crossings. Notably, there exists a counter-intuitive trade-off between #CR and WL, minimizing #CR may lead to long detours that increase propagation loss, resulting in a higher overall . Except for the Light_b case, LiDAR achieves the lowest across all other cases while maintaining crossing-optimal (#CR=0), DRV-free layouts.
@@ VI-B Results on Compact Benchmarks
We further evaluate the routing performance on compact benchmarks characterized by higher component density and more topological crossings. For these cases, only a single round of rip-up and reroute is applied, as limited layout space makes additional iterations ineffective. Figure 15 shows the final routed layouts generated by LiDAR 2.0.
As shown in Table VI, LiDAR 2.0 achieves nearly DRV-free results across all benchmarks, with an average 16% reduction in insertion loss and a 5.16 speedup over PROTON, and a 6.95 speedup and 9% lower insertion loss compared to LiDAR 1.0. On the TeMPO benchmark, PROTON shows a much lower runtime due to complete port blocking and early termination without successful routing.
On the compact cases, LiDAR 1.0 also shows increased DRVs and runtime, especially on newly added cases. In contrast, LiDAR 2.0 adopts a hierarchical routing strategy that enables subcircuit reuse in designs such as TeMPO and Bennes, significantly improving runtime. Building on this, intra-group routing order optimization and proactive crossing space preservation further reduce conflicts in multiport components, improve solution quality, and avoid unnecessary search. As a result, LiDAR 2.0 successfully achieves DRV-free routing even on high-crossing-density Bennes benchmarks. For GWOR, which lacks group information, LiDAR 2.0 still generates DRV-free results by dynamically refining the routing order and port regions during rip-up and reroute.
Figure 16 presents the runtime breakdown across various routing stages on Bennes_3232. The substantial speedup achieved by LiDAR 2.0 stems from its ability to minimize routing conflicts, eliminate redundant searches and RR iterations, and hierarchically reuse routing results.
@@ VI-C Results on Different Crossing Sizes and Bend Radii
| Bennes_1616_C | 23 | 3.57 | 17.15 | 0 | 33 | 23 | 3.45 | 17.13 | 0 | 34 | 23 | 3.34 | 17.15 | 0 | 34 |
| Bennes_1616_L | 22 | 3.70 | 16.68 | 0 | 34 | 22 | 3.57 | 16.62 | 0 | 28 | 23 | 3.49 | 17.23 | 0 | 35 |
| Metrics | GWOR_3232 | Clements_1616_C | ADEPT_1616_C | Bennes_1616_C | TeMPO_1616_C | ||||||||
@@
| Time (s) | 10 | 10 | 10 | 14 | 16 | 96 | 111 | 298 | 339 | 33 | 33 | 44 | 172 |
Different PIC technologies impose varying constraints on crossing sizes and bending radii. Table VII and Table VIII show the routing results of LiDAR 2.0 under different crossing sizes and bend radii. Benchmarks with an L suffix indicate larger component pitch. As expected, larger crossing sizes and bend radii increase area usage and routing difficulty. Still, LiDAR 2.0 consistently generates DRV-free layouts across compact and spacious benchmarks, with only minor DRVs in ADEPT cases. LiDAR 2.0 supports the insertion of various crossing sizes and remains robust under different bend radii. It successfully routes GWOR_3232 even with large bend radii up to 80 , which is typical in silicon nitride platforms. For silicon photonics, where the bending radius is typically 5–10 , LiDAR 2.0 performs reliably. Feasible solutions remain available at 15 despite minor DRVs, except for the TeMPO and Bennes benchmarks, where no valid routing solution exists due to insufficient spacing.
@@ VI-D Discussion
Crossing-Waveguide Optimization. As shown in Table IV (PROTON [19] vs. LiDAR 2.0), our proposed crossing optimization strategy introduces an extra runtime penalty on the Light benchmark, but it leads to higher solution quality as it reduces the crossings between IOs and the switch given the high crossing penalty.
Port-Group-based Congestion Penalty (GCP). We evaluate the effectiveness of our proposed group-based congestion penalty in guiding crossing optimization under different crossing insertion loss settings, as shown in Table IX. Specifically, we compare scenarios with high crossing IL () and low crossing IL (). When the crossing IL is high, our method effectively avoids crossings to minimize the maximum insertion loss (). In contrast, under low crossing IL, it prefers shorter waveguide paths with more crossings to further optimize .
Without the group-based penalty, however, the router tends to use more crossings as increases. This is because, under congestion from previously routed nets, it struggles to find low-crossing paths. By introducing the group-based penalty, routing conflicts are significantly reduced, enabling more efficient decisions with fewer crossings and lower overall IL. As illustrated in Fig. 17, the coefficient functions as a flexible control parameter, allowing users to balance between waveguide length and the number of crossings based on specific PIC performance requirements.
Offset Neighbor. Offset neighbors help eliminate unnecessary bending detours caused by port misalignment in PICs, thereby reducing both routing conflicts and search overhead. To further demonstrate their effectiveness, we evaluate the router on a compact TeMPO_C+ variant with x/y pitch reduced from 25/50 to 20/15 , as shown in Table X. In such highly constrained layouts, detours without offset neighbors often result in design rule violations. Even in more spacious cases, enabling offset neighbors consistently leads to faster runtime compared to routing without them.
@@ VII Conclusion
In this work, we first introduce LiDAR, an open-source automated detailed routing tool specifically designed for photonic integrated circuits. LiDAR features a non-Manhattan curvy-aware A∗ search engine with accessibility-enhanced port assignment, adaptive crossing insertion, congestion-aware group-based net ordering and objective, and crossing optimization scheme to handle unique PIC routing constraints while optimizing critical path insertion loss. To further alleviate routing conflicts and improve efficiency, we extend LiDAR into a more scalable and conflict-resilient framework, LiDAR 2.0, by introducing hierarchical routing reuse, intra-group net ordering, offset-neighbor search, and crossing-space preservation. These enhancements substantially improve routability under tight layout constraints and high crossing density. LiDAR 2.0 achieves nearly DRV-free layouts across challenging benchmarks, delivering up to 9% lower insertion loss and 6.95 speedup compared to LiDAR 1.0, advancing the automation frontier for large-scale PIC physical design.

### v2

@@ VI Evaluation Results
The proposed photonic detailed routing framework is implemented in Python based on GDSFactory 8 [31] libraries. All evaluations are conducted on a Linux server with a 128-core AMD EPYC 7763 CPU and 512-GB memory. Note that runtime and results may slightly differ from the ISPD version due to platform and GDSFactory version variations.
@@
A∗ Search Parameters. Referring to the parameters in Table V and scaling them to a unified micrometer unit, the base propagation, bending, and crossing weights are set to 1.5, 50, and 5000, respectively. On top of these, we further adjust the weights according to the characteristics of each benchmark. For the photonic tensor core, where crossings are inherent and unavoidable, we set the crossing weight equal to the bending weight to speed up search. For the optical switch, the crossing weight remains 5000, while the bending weight is also set to 5000 to reduce bends; in long waveguides with large radii, shorter wirelength can offset added bend loss. The group congestion penalty (GCP) parameter is determined by the detour distance, as saving routing resources through GCP may conflict with wirelength minimization. We set this value to 500, corresponding to reserving approximately 200 m of spacing. The crossing spacing coefficient is set similarly, but its checking range is limited to a single crossing.
@@ VI-A Results on Spacious Benchmarks
Analysis of PTC Results. Despite the Clements circuit having a regular topology with no inherent crossings and relatively spacious layouts, insufficient planning and lack of reserved access space lead to unnecessary crossings and DRVs in PROTON. These routing conflicts become more severe in ADEPT benchmarks, which feature dense multi-port MMIs and a large number of crossings. By incorporating Accessibility-Enhanced Port Assignment and Port-Group-Based Net Ordering, LiDAR 2.0 effectively mitigates these conflicts and reduces unnecessary rip-up and reroute. As circuit scale increases, PROTON suffers from escalating #DRVs and runtime, while LiDAR 2.0 consistently produces DRV-free, low-IL layouts.
Analysis of Optical Switch Results. Light benchmark provides abundant routing resources and is primarily used to evaluate the router’s ability to optimize insertion loss and crossings. Notably, there exists a counter-intuitive trade-off between #CR and WL; minimizing #CR may lead to long detours that increase propagation loss, resulting in a higher overall . Except for the Light_b case, LiDAR achieves the lowest across all other cases while maintaining crossing-optimal (#CR=0), DRV-free layouts.
@@ VI-B Results on Compact Benchmarks
We further evaluate routing on compact benchmarks with higher component density and more crossings. Only one rip-up and reroute round is applied, as limited space makes further iterations ineffective. Figure 15 shows the final layouts by LiDAR 2.0.
As in Table VI, LiDAR 2.0 delivers nearly DRV-free results with an average 16% IL reduction, 5.16 speedup over PROTON, and 6.95 speedup with 9% lower IL vs. LiDAR 1.0. On TeMPO, PROTON runs faster only because complete port blocking forces early termination. For compact cases, LiDAR 1.0 shows more DRVs and runtime, especially on new benchmarks. In contrast, LiDAR 2.0’s hierarchical routing reuses subcircuits (TeMPO, Benes), greatly improving runtime. Intra-group ordering and proactive crossing-space reservation further reduce conflicts, enhance quality, and avoid redundant search. Thus, LiDAR 2.0 achieves DRV-free routing even on high-crossing Benes. For GWOR, lacking group data, LiDAR 2.0 still succeeds by dynamically refining order and port regions during reroute.
Figure 16 shows the runtime breakdown on Benes_3232. The speedup of LiDAR 2.0 comes from reducing conflicts, eliminating redundant search and RR iterations, and hierarchically reusing results.
@@ VI-C Results on Different Crossing Sizes and Bend Radii
| Benes_1616_C | 23 | 3.57 | 17.15 | 0 | 33 | 23 | 3.45 | 17.13 | 0 | 34 | 23 | 3.34 | 17.15 | 0 | 34 |
| Benes_1616_L | 22 | 3.70 | 16.68 | 0 | 34 | 22 | 3.57 | 16.62 | 0 | 28 | 23 | 3.49 | 17.23 | 0 | 35 |
| Metrics | GWOR_3232 | Clements_1616_C | ADEPT_1616_C | Benes_1616_C | TeMPO_1616_C | ||||||||
@@
| Time (s) | 10 | 10 | 10 | 14 | 16 | 96 | 111 | 298 | 339 | 33 | 33 | 169 | 172 |
Different PIC technologies impose varying constraints on crossing sizes and bending radii. Table VII and Table VIII show the routing results of LiDAR 2.0 under different crossing sizes and bend radii. We synthetically stretch the device only to increase the routing difficulty and test our router’s capability, regardless of its actual IL value. Benchmarks with an L suffix indicate larger component pitch. As expected, larger crossing sizes and bend radii increase area usage and routing difficulty. Still, LiDAR 2.0 consistently generates DRV-free layouts across compact and spacious benchmarks, with only minor DRVs in ADEPT cases. LiDAR 2.0 supports the insertion of various crossing sizes and remains robust under different bend radii. It successfully routes GWOR_3232 even with large bend radii up to 80 , which is typical in silicon nitride platforms. For silicon photonics, where the bending radius is typically 5–10 , LiDAR 2.0 performs reliably. Feasible solutions remain available at 15 despite minor DRVs, except for the TeMPO and Benes benchmarks, where no valid routing solution exists due to insufficient spacing.
@@ VI-D Discussion
| Metrics | Light_a | ADEPT_44 | ||
| Manual | LiDAR 2.0 | Manual | LiDAR 2.0 | |
| #CR | 0 | 0 | 10 | 10 |
| WL (mm) | 31.89 | 29.99 | 2.30 | 2.33 |
| 7.85 | 7.61 | 12.95 | 13.14 | |
| DRV | 0 | 0 | 0 | 0 |
| Time | 2h | 103s | 2h | 26s |
@@
| Benchmarks | w/o Staggered offset | w/o crossing space | w/o group-based net order | |||
| #DRV | #DRV | #DRV | ||||
| ADEPT_1616_C | 18.95 | 37 | 19.97 | 6 | 21.63 | 4 |
| ADEPT_1616_L | 19.13 | 33 | 20.34 | 3 | 21.95 | 3 |
| TeMPO_1616_C | 28.49 | 27 | 28.31 | 0 | 31.87 | 16 |
| TeMPO_1616_L | 28.72 | 26 | 28.22 | 0 | 32.64 | 16 |
| Benes_1616_C | 14.62 | 21 | 16.24 | 32 | 16.83 | 2 |
| Benes_1616_L | 14.37 | 21 | 16.93 | 30 | 17.21 | 4 |
@@
Manual Routing Result. We conducted a manual routing comparison on small circuits, Light_a and ADEPT_44, as shown in TABLE IX. It can be observed that the routing quality of LiDAR 2.0 is comparable to that of manual routing, and even achieves lower insertion loss on Light_a. However, manual routing required approximately two hours to complete for these small circuits, and as the circuit scale increases, manual efforts become impractical.
Spacing Reservation. We propose several techniques for reserving spacing, including staggered offset port region, reserved crossing spacing, group congestion penalty, and group-based net ordering to mitigate routing conflicts in multiport components. TABLE XI highlights their impact on representative circuits. It can be observed that the staggered offset port region is most critical: removing it causes many routing failures. Without the crossing space penalty, the Benes circuit suffers heavily due to consecutive crossings, while TEMPO is less affected since its crossings are independent. Disabling group-based net ordering, however, triggers numerous DRVs in TEMPO because of high-fanout components.
We also assess the proposed GCP in guiding crossing optimization under different crossing weights (Table X). With larger , the router avoids crossings to reduce maximum insertion loss (); with smaller , it favors shorter paths even with extra crossings, further improving .
Without GCP, more crossings arise as increases, since earlier nets congest low-crossing paths. With GCP, routing conflicts drop notably, allowing efficient paths with fewer crossings and lower IL. In Fig. 17, flexibly balances waveguide length and crossings to meet performance needs.
Crossing-Waveguide Optimization. As shown in Table IV (PROTON [12] vs. LiDAR 2.0), our proposed crossing optimization strategy introduces an extra runtime penalty on the Light benchmark, but it leads to higher solution quality as it reduces the crossings between IOs and the switch, given the high crossing penalty.
Offset Neighbor. Offset neighbors help eliminate unnecessary bending detours caused by port misalignment, thereby reducing routing conflicts and search overhead. We evaluate them on a compact TeMPO_C+ variant with x/y pitch reduced from 25/50 to 20/15 (Table XII). In such constrained layouts, routing without offset neighbors often causes rule violations. Even in spacious cases, enabling them yields consistently faster runtime.
@@ VII Conclusion
In this work, we present LiDAR, an open-source detailed router for PICs. It features a non-Manhattan curvy-aware A∗ engine with enhanced port assignment, adaptive crossing insertion, congestion-aware group-based net ordering, and crossing optimization to address unique PIC constraints while minimizing critical path insertion loss. We further extend it into a scalable, conflict-resilient framework, LiDAR 2.0, by adding hierarchical routing reuse, intra-group ordering, offset-neighbor search, and crossing-space preservation. These enhancements markedly improve routability under tight layouts and high crossing density. LiDAR 2.0 achieves nearly DRV-free layouts on challenging benchmarks, with up to 9% lower IL and 6.95 speedup over LiDAR 1.0, advancing automation for large-scale PIC design.